### v2

# Compression and complexity for sumset sizes in additive number theory

Abstract.
The study of sums of finite sets of integers has mostly concentrated on sets with small sumsets (Freiman’s theorem and related work) and on sets with large sumsets (Sidon sets and $B_{h}$-sets). This paper considers the sets ${\mathcal{R}}_{\mathbf{Z}}(h,k)$ and ${\mathcal{R}}_{{\mathbf{Z}}^{n}}(h,k)$ of all sizes of $h$-fold sums of sets of $k$ integers or of $k$ lattice points, and the geometric and computational complexity of the sets ${\mathcal{R}}_{\mathbf{Z}}(h,k)$ and ${\mathcal{R}}_{{\mathbf{Z}}^{n}}(h,k)$. For sumsets $hA$ with large diameter, there is a compression algorithm to construct sets $A^{\prime}$ with $|hA^{\prime}|=|hA|$ and small diameter.
Key words and phrases:
Sumset, restricted sumset, distribution of of sumset sizes, geometric compression, computational complexity, additive number theory, combinatorial number theory
2000 Mathematics Subject Classification:
11B75, 11B05, 11B13, 11B30, 11P70, 11Y16, 11Y55
Supported in part by PSC-CUNY Research Award Program grant 66197-00 54.
1. Geometric and computational complexity
Let $X$ be an additive abelian semigroup and let $A$ be a nonempty subset of $X$. The $h$-fold sum of $A$ is the set of all sums of $h$ not necessarily distinct elements of $A$. We study the sizes of $h$-fold sumsets of $k$-element subsets of $X$:
$\mathcal{R}_{X}(h,k)=\{|hA|:A\subseteq X\text{ and }|A|=k\}.$
A central problem in additive number theory is to compute and understand the sumset size set $\mathcal{R}_{X}(h,k)$. This is unsolved even in the classical cases when $X$ is the semigroup $\mathbf{N}_{0}$ of nonnegative integers, the groups $\mathbf{Z}$, $\mathbf{Q}$, and $\mathbf{R}$ of integers, rational numbers, and real numbers, and the groups $\mathbf{Z}^{n}$, $\mathbf{Q}^{n}$, and $\mathbf{R}^{n}$ of lattice points with integral, rational, and real coordinates. If sets $A$ an $A^{\prime}$ are affinely equivalent, that is, if
$A^{\prime}=\lambda\ast A+\mu=\{\lambda a+\mu:a\in A\}$
for real numbers $\lambda\neq 0$ and $\mu$, then $|hA^{\prime}|=|hA|$. It follows that, for all positive integers $h$ and $k$,
(1) $\mathcal{R}_{\mathbf{N}_{0}}(h,k)=\mathcal{R}_{\mathbf{Z}}(h,k)=\mathcal{R}_{\mathbf{Q}}(h,k)$
and
(2) $\mathcal{R}_{\mathbf{N}_{0}^{n}}(h,k)=\mathcal{R}_{\mathbf{Z}^{n}}(h,k)=\mathcal{R}_{\mathbf{Q}^{n}}(h,k).$
We have
(3) $\min\mathcal{R}_{\mathbf{Z}}(h,k)=\min\mathcal{R}_{\mathbf{R}}(h,k)=hk-h+1$
and
(4) $\max\mathcal{R}_{\mathbf{Z}}(h,k)=\max\mathcal{R}_{\mathbf{R}}(h,k)=\binom{h+k-1}{k-1}$.
In Theorem 5, we prove that
$\mathcal{R}_{\mathbf{Z}}(h,k)=\mathcal{R}_{\mathbf{Z}^{n}}(h,k)$
for all positive integers $n$. Of course,
$\mathcal{R}_{\mathbf{Z}}(h,k)\subseteq\mathcal{R}_{\mathbf{R}}(h,k)$
but it is not known if $\mathcal{R}_{\mathbf{Z}}(h,k)=\mathcal{R}_{\mathbf{R}}(h,k)$ for some or all positive integers $h$ and $k$.
This paper considers the geometric and computational complexity of the sets $\mathcal{R}_{\mathbf{Z}}(h,k)$ and $\mathcal{R}_{\mathbf{Z}^{n}}(h,k)$. This problem was introduced in [8, 9, 10] and basic results can be found there. For all positive integers $h$ and $k$, we have
$\mathcal{R}_{\mathbf{Z}}(h,1)=\{1\},\qquad\mathcal{R}_{\mathbf{Z}}(h,2)=\{h+1\},\qquad\mathcal{R}_{\mathbf{Z}}(1,k)=\{k\}$
and, from [9],
$\mathcal{R}_{\mathbf{Z}}(2,k)=\left[2k-1,\binom{k+1}{2}\right$]
where, for all $u,v\in\mathbf{R}$,there is the integer interval
$u,v]=\{n\in\mathbf{Z}:u\leq n\leq v\}.$
Problem 1.
The fundamental problem is to compute $\mathcal{R}_{\mathbf{Z}}(h,k)$ for all $h\geq 3$ and $k\geq 3$.
The diameter of a nonempty subset $A$ of a metric space is $\operatorname{\text{diam}}(A)=\sup\{d(x,y):x,y\in A\}$. Let $A=\{a_{1},a_{2},\ldots,a_{k-1},a_{k}\}$ be a set of $k$ nonnegative integers with $0=a_{1}<a_{2}<\cdots<a_{k-1}<a_{k}$ and $|hA|=t\in\mathcal{R}_{\mathbf{Z}}(h,k)$. If $a_{k}>ha_{k-1}+1$, then the “compressed” set $A^{\prime}=\{a_{1},a_{2},\ldots,a_{k-1},a^{\prime}_{k}\}$ with $a^{\prime}_{k}=ha_{k-1}+1$ also satisfies $|A^{\prime}|=k$ and $|hA^{\prime}|=t$ (Theorem 2), and
$\operatorname{\text{diam}}(A^{\prime})=ha_{k-1}+1<a_{k}=\operatorname{\text{diam}}(A).$
For example, if $A=\{0,1,3,20,200\}$, then $|A|=5$ and $|4A|=65$. If $A^{\prime}=\{0,1,3,20,81\}$, then $|A^{\prime}|=5$ and $|4A^{\prime}|=65$. Indeed, for all $b\geq 81$, if $A^{\prime}=\{0,1,3,20,b\}$, then $|A^{\prime}|=5$ and $|4A^{\prime}|=65$. However, if $A^{\prime}=\{0,1,3,20,80\}$, then $|A^{\prime}|=5$ and $|4A^{\prime}|=64$.
In this example, we reduced the diameter of the set $A$ by compressing its largest element. Consider reducing the diameter of $A$ by compressing the gap between “interior” elements of $A$. For example, the set $A^{\prime}=\{0,1,3,20,81\}$ satisfies $|A^{\prime}|=5$ and $|4A^{\prime}|=65$. Shrinking the gap between the elements 3 and 20, we see that the set $A^{\prime\prime}=\{0,1,3,20-c,81-c\}$ satisfies $|A^{\prime\prime}|=5$ and $|4A^{\prime\prime}|=65$ for all $c\in[0,7]$, but, with $c=8$ and $A^{\prime\prime}=\{0,1,3,12,73\}$, we obtain $|4A^{\prime\prime}|=64$.
Problem 2.
The geometric complexity problem is to describe, for positive integers $h$ and $k$, the shapes of “compressed” sets of $k$ integers or lattice points with given $h$-fold sumset size $t$.
There are also computational complexity problems for sumsets. By (3) and (4), the set $\mathcal{R}_{\mathbf{Z}}(h,k)$ is finite. For all $t\in\mathcal{R}_{\mathbf{Z}}(h,k)$, there is an integer $N_{t}$ and a set $A\subseteq[0,N_{t}-1]$ with $|A|=k$ and $|hA|=t$. Let $N=\max\{N_{t}:t\in\mathcal{R}_{\mathbf{Z}}(h,k)\}$. Then
$\mathcal{R}_{\mathbf{Z}}(h,k)=\{|hA|:A\subseteq[0,N-1]\text{ and }|A|=k\}.$
Let $N(h,k)$ be the least integer $N$ that satisfies this condition. For all $h\geq 1$ and $k\geq 1$ we have
$N(h,1)=1,\qquad N(1,k)=k,\operatorname{\qquad\text{and}\qquad}N(h,2)=2.$
Nathanson [9] proved that
$N(2,k)\leq 2^{k}\operatorname{\qquad\text{and}\qquad}N(3,3)=5.$
Theorem 8 gives the upper bound
$N(h,k)<4(8h)^{k-1}$
for all $h\geq 3$ and $k\geq 3$.
Problem 3.
The computational complexity problem is to determine $N(h,k)$ for all $h\geq 3$ and $k\geq 3$.
For related work on sumset sizes, see [2, 5, 6, 8, 12].
2. Compression theorem for integers
By abuse of notation, we write
$A=\{a_{1}<a_{2}<\cdots<a_{k}\}$
if $A=\{a_{1},a_{2},\ldots,a_{k}\}$ is a set of $k$ integers and $a_{1}<a_{2}<\cdots<a_{k}$.
Theorem 1.
Let $k\geq 3$ and let $A=\{a_{1}<a_{2}<\cdots<a_{k}\}$ be a set of $k$ integers. Let $h$ be a positive integer. If there exist an integer $j\in[1,k-1]$ and a positive integer $\delta_{j}$ such that
(5) $a_{j+1}-a_{j}=1+\delta_{j}+(h-1)\max(a_{j}-a_{1},a_{k}-a_{j+1})$
then the integer $j$ is unique and the strictly increasing sequence of $k$ integers defined by
[…]
and the set $A^{\prime}=\{a^{\prime}_{1}<a^{\prime}_{2}<\cdots<a^{\prime}_{k}\}$ satisfy
(7) $|hA^{\prime}|=|hA|.$
Moreover,
(8) $a^{\prime}_{j+1}-a^{\prime}_{j}=1+(h-1)\max(a^{\prime}_{j}-a^{\prime}_{1},a^{\prime}_{k}-a^{\prime}_{j+1})$
and
(9) $\operatorname{\text{diam}}(A^{\prime})=\operatorname{\text{diam}}(A)-\delta_{j}\leq\operatorname{\text{diam}}(A)-1.$
Proof.
Suppose there exist distinct integers $j_{1}\in[1,k-1]$ and $j_{2}\in[1,k-1]$ that satisfy (5). If $j_{1}<j_{2}$, then
$a_{j_{1}+1}-a_{j_{1}}$ $>(h-1)\max(a_{j_{1}}-a_{1},a_{k}-a_{j_{1}+1})$
$\geq(h-1)(a_{k}-a_{j_{1}+1})$
$\geq a_{j_{2}+1}-a_{j_{2}}$
$>(h-1)\max(a_{j_{2}}-a_{1},a_{k}-a_{j_{2}+1})$
$\geq(h-1)(a_{j_{2}}-a_{1})$
$\geq a_{j_{1}+1}-a_{j_{1}}$
which is absurd. Therefore, at most one integer $j\in[1,k-1]$ satisfies (5).
Condition (5) implies
$(h-1-i)$ $(a_{j}-a_{1})+i(a_{k}-a_{j+1})$
(10) $\leq(h-1)\max(a_{j}-a_{1},a_{k}-a_{j+1})$
$<a_{j+1}-a_{j}-\delta_{j}$
for all $i\in[0,h-1]$.
Let
$A=B\cup C$
where
$B=\{a_{1}<\cdots<a_{j}\}$
and
$C=\{a_{j+1}<\cdots<a_{k}\}.$
The sumset $hA$ decomposes as follows:
$hA=h(B\cup C)=\bigcup_{i=0}^{h}\left((h-i)B+iC\right).$
We have
$(h-i)B+iC$ $\subseteq[(h-i)a_{1},(h-i)a_{j}]+[ia_{j+1},ia_{k}]$
$=[(h-i)a_{1}+ia_{j+1},(h-i)a_{j}+ia_{k}].$
For all $i\in[0,h-1]$,
$\max\left((h-i)B+iC\right)=(h-i)a_{j}+ia_{k}$
and
$\min\left((h-(i+1))B+(i+1)C\right)=(h-1-i)a_{1}+(i+1)a_{j+1}.$
The $h+1$ sets (hi)B+iC(h-i)B+iC are pairwise disjoint if, for all $i\in[0,h-1]$,
$(h-i)a_{j}+ia_{k}<(h-1-i)a_{1}+(i+1)a_{j+1}$
or, equivalently, if
$(h-1-i)(a_{j}-a_{1})+i(a_{k}-a_{j+1})<a_{j+1}-a_{j}.$
This is inequality (2) and so the $h+1$ sets (hi)B+iC(h-i)B+iC are pairwise disjoint and
(11) $|hA|=\sum_{i=0}^{h}\left|(h-i)B+iC\right|.$
We have
$a^{\prime}_{j}-a^{\prime}_{1}$ $=a_{j}-a_{1}$
$a^{\prime}_{k}-a^{\prime}_{j+1}$ $=(a_{k}-\delta_{j})-(a_{j+1}-\delta_{j})=a_{k}-a_{j+1}$
and, for all $\ell\in[1,k-1]\setminus\{j\}$,
$a^{\prime}_{{\ell}+1}-a^{\prime}_{\ell}$ $=a_{{\ell}+1}-a_{\ell}\geq 1.$
Condition (5) gives
$a^{\prime}_{j+1}-a^{\prime}_{j}$ $=(a_{j+1}-\delta_{j})-a_{j}$
$=1+(h-1)\max(a_{j}-a_{1},a_{k}-a_{j+1})$
$=1+(h-1)\max(a^{\prime}_{j}-a^{\prime}_{1},a^{\prime}_{k}-a^{\prime}_{j+1}).$
This proves (8). Moreover,
$a^{\prime}_{1}<\cdots<a^{\prime}_{j-1}<a^{\prime}_{j}<a^{\prime}_{j+1}<\cdots<a^{\prime}_{k}$
and the set
$A^{\prime}=\{a^{\prime}_{1}<a^{\prime}_{2}<\cdots<a^{\prime}_{k}\}$
satisfies
$\operatorname{\text{diam}}(A^{\prime})=a^{\prime}_{k}-a^{\prime}_{1}=a_{k}-a_{1}-\delta_{j}=\operatorname{\text{diam}}(A)-\delta_{j}\leq\operatorname{\text{diam}}(A)-1.$
This proves (9).
We shall prove that $|hA^{\prime}|=|hA|$. Let
$A^{\prime}=B^{\prime}\cup C^{\prime}$
where
$B^{\prime}=\{a^{\prime}_{1}<\cdots<a^{\prime}_{j}\}=\{a_{1}<\cdots<a_{j}\}=B$
and
$C^{\prime}=\{a^{\prime}_{j+1}<\cdots<a^{\prime}_{k}\}=\{a_{j+1}-\delta_{j}<\cdots<a_{k}-\delta_{j}\}=C-\delta_{j}.$
Then
$hA^{\prime}=h(B^{\prime}\cup C^{\prime})=\bigcup_{i=0}^{h}\left((h-i)B^{\prime}+iC^{\prime}\right).$
We have
$(h-i)B^{\prime}+iC^{\prime}=(h-i)B+i(C-\delta_{j})=(h-i)B+iC-i\delta_{j}$
and so
$\left|(h-i)B^{\prime}+iC^{\prime}\right|=\left|(h-i)B+iC\right|.$
If the sets $(h-i)B^{\prime}+iC^{\prime}$ are pairwise disjoint for all $i\in[0,h]$, then
$|hA^{\prime}|=\sum_{i=0}^{h}\left|(h-i)B^{\prime}+iC^{\prime}\right|=\sum_{i=0}^{h}\left|(h-i)B+iC\right|=|hA|.$
Observe that
$(h-i)B^{\prime}+iC^{\prime}$ $\subseteq[(h-i)a_{1},(h-i)a_{j}]+[ia_{j+1}-i\delta_{j},ia_{k}-i\delta_{j}]$
$=[(h-i)a_{1}+ia_{j+1}-i\delta_{j},(h-i)a_{j}+ia_{k}-i\delta_{j}].$
For all $i\in[0,h-1]$, the sets $(h-i)B^{\prime}+iC^{\prime}$ and $(h-i-1)B^{\prime}+(i+1)C^{\prime}$ are pairwise disjoint if
$(h-i)a_{j}+ia_{k}-i\delta_{j}<(h-i-1)a_{1}+(i+1)a_{j+1}-(i+1)\delta_{j}$
or, equivalently, if
$(h-i-1)(a_{j}-a_{1})+i(a_{k}-a_{j+1})<a_{j+1}-a_{j}-\delta_{j}.$
This is exactly inequality (2) and so $|hA^{\prime}|=|hA|$. This completes the proof. ∎
Theorem 2.
Let $A=\{a_{1}<\cdots<a_{k-1}<a_{k}\}$ be a set of $k$ integers with $a_{1}=0$. For all positive integers $h$, if $a_{k}>1+ha_{k-1}$, then the set $A^{\prime}=\{a_{1}<\cdots<a_{k-1}<a^{\prime}_{k}\}$ with $a^{\prime}_{k}=1+ha_{k-1}$ satisfies $|hA^{\prime}|=|hA|$.
Proof.
If $a_{k}>1+ha_{k-1}$, then there is a positive integer $\delta_{k-1}$ such that
$a_{k}-a_{k-1}$ $=1+\delta_{k-1}+(h-1)a_{k-1}$
$=1+\delta_{k-1}+(h-1)\max(a_{k-1}-1,a_{k}-a_{k}).$
The construction in Theorem 1 produces the set $A^{\prime}=\{a_{1}<\cdots<a_{k-1}<a^{\prime}_{k}\}$ with
$a^{\prime}_{k}=a_{k}-\delta_{k-1}=1+ha_{k}$
and $|hA^{\prime}|=|hA|$. This completes the proof. ∎
Theorem 3.
Let $k\geq 3$ and let $A=\{a_{1}<a_{2}<\cdots<a_{k}\}$ be a set of $k$ integers. For every positive integer $h$, there is a set $A^{\prime}=\{a^{\prime}_{1}<a^{\prime}_{2}<\cdots<a^{\prime}_{k}\}$ such that $|hA^{\prime}|=|hA|$ and
(12) $a^{\prime}_{j+1}-a^{\prime}_{j}\leq 1+(h-1)\max(a^{\prime}_{j}-a^{\prime}_{1},a^{\prime}_{k}-a^{\prime}_{j+1})$
for all $j\in[1,k-1]$.
Proof.
If the set $A$ does not satisfy condition (5) for some $j\in[1,k-1]$, then apply Theorem 1 and obtain a set $A^{\prime}$ with $|hA^{\prime}|=|hA|$ and $\operatorname{\text{diam}}(A^{\prime})<\operatorname{\text{diam}}(A)$. Repeat this argument if the set $A^{\prime}$ does not satisfy inequality (12) for some $j\in[1,k-1]$. Continuing this procedure, we obtain a sequence of sets whose diameters form a strictly decreasing sequence of positive integers, and so the process must stop after finitely many iterations. This completes the proof. ∎
Theorem 2 is a special case of the following result.
Theorem 4.
Let $k\geq 3$ and let $A=\{a_{1}<a_{2}<\cdots<a_{k}\}$ be a set of $k$ integers such that $a_{1}=0$ and
(13) $a_{j+1}-a_{j}\leq 1+(h-1)\max(a_{j},a_{k}-a_{j+1})$
for all $j\in[1,h-1]$, There is a unique integer $r$ in [2,k1][2,k-1] such that
$a_{r-1}+a_{r}<a_{k}\leq a_{r}+a_{r+1}.$
For all $j\in[r,k-1]$,
$a_{k}\leq 1+h+h^{2}+\cdots+h^{k-j-1}+h^{k-j}a_{j}.$
Proof.
Because $a_{1}=0$ and $k\geq 3$, we have
$a_{1}+a_{2}=a_{2}<a_{k}<a_{k-1}+a_{k}.$
Choose $r\in[2,k-1]$ such that
$a_{r-1}+a_{r}<a_{k}\leq a_{r}+a_{r+1}.$
For all $j\in[r,k-1]$, we have
$a_{k}\leq a_{r}+a_{r+1}\leq a_{j}+a_{j+1}$
and so
$a_{k}-a_{j+1}\leq a_{j}.$
It follows that
$a_{j+1}-a_{j}\leq 1+(h-1)\max(a_{j},a_{k}-a_{j+1})=1+(h-1)a_{j}$
and
$a_{j+1}\leq 1+ha_{j}.$
If $r\leq j\leq j+1\leq k-1$, then
$a_{j+2}$ $\leq 1+ha_{j+1}$
$\leq 1+h(1+ha_{j})$
$=1+h+h^{2}a_{j}.$
By induction, if $r\leq j\leq j+\ell\leq k$, then
$a_{j+\ell}\leq 1+h+h^{2}+\cdots+h^{\ell-1}+h^{\ell}a_{j}.$
Choosing $\ell=k-j$, we obtain
$a_{k}\leq 1+h+h^{2}+\cdots+h^{k-j-1}+h^{k-j}a_{j}.$
This completes the proof. ∎
For example, let $h=3$ and $k=7$. Consider the set
$A=\{0,2,7,11,70,85,91\}.$
We have $k=|A|=7$ and $|hA|=|3A|=80$. Applying Theorem 3, we obtain
$59=70-11=1+\delta_{4}+2\max(11,91-70)=\delta_{4}+43$
and so $\delta_{4}=16$ and
$A^{\prime}=\{0,2,7,11,54,69,75\}.$
Note that
$43=54-11=1+2\max(11,75-54).$
The sumsets $3A$ and $3A^{\prime}$ are different, that is, $3A\neq 3A^{\prime}$, but
$|3A|=|3A^{\prime}|=80.$
The compressed set $A^{\prime}$ satisfies condition (13) for all $j\in[1,6]$. We have
$11+54=65<75<123=54+69.$
Applying Theorem 4 with $r=5$, we obtain
$75$ $<1+3+3^{2}\cdot 54$
$75$ $<1+3\cdot 69.$
3. Additive isomorphism $\mathcal{R}_{\mathbf{Z}^{n}}(h,k)=\mathcal{R}_{\mathbf{Z}}(h,k)$
Let $X$ and $Y$ be additive abelian groups. Let $A$ be a finite subset of $X$ and $B$ a finite subset of $Y$. The function $f:A\rightarrow B$ is a Freiman isomorphism of order $h$ if $f$ is a bijection and, for all $a_{1},\ldots,a_{h},a^{\prime}_{1},\ldots,a^{\prime}_{h}\in A$, we have
$a_{1}+\cdots+a_{h}=a^{\prime}_{1}+\cdots+a^{\prime}_{h}$
if and only if
$f(a_{1})+\cdots+f(a_{h})=f(a^{\prime}_{1})+\cdots+f(a^{\prime}_{h}).$
Thus, $f$ induces a bijection from $hA$ to $hB$ and so $|hA|=|hB|$. For applications of Freiman isomorphisms, see [7].
The $\ell^{\infty}$-norm of a vector $\mathbf{x}=(x_{1},\ldots,x_{n})\in\mathbf{R}^{n}$ is
$\|\mathbf{x}\|_{\infty}=\max(|x_{j}|:j\in[1,n]\}.$
The proof of the following result is a simple application of Freiman isomorphisms.
Theorem 5.
For all positive integers $h$, $k$, and $n$,
$\mathcal{R}_{\mathbf{Z}^{n}}(h,k)=\mathcal{R}_{\mathbf{Z}}(h,k).$
Proof.
The group $\mathbf{Z}^{n}$ contains subgroups isomorphic to $\mathbf{Z}$, and so $\mathcal{R}_{\mathbf{Z}}(h,k)\subseteq\mathcal{R}_{\mathbf{Z}^{n}}(h,k)$. To prove that $\mathcal{R}_{\mathbf{Z}^{n}}(h,k)\subseteq\mathcal{R}_{\mathbf{Z}}(h,k)$, it suffices to show that, for every positive integer $h$ and every finite set $A$ of lattice points, there is a finite set $B$ of integers that is Freiman isomorphic of order $h$ to $A$.
Let $A\subseteq\mathbf{Z}^{n}$ with $|A|=k$. By translation, we can assume that $A\subseteq\mathbf{N}_{0}^{n}$. Choose a positive integer
$g>h\max\{\|\mathbf{a}\|_{\infty}:\mathbf{a}\in A\}.$
Define $f:\mathbf{Z}^{n}\rightarrow\mathbf{Z}$ by
$f(a_{1},\ldots,a_{n})=\sum_{j=1}^{n}a_{j}g^{j-1}$
and let
$B=f(A)\subseteq\mathbf{N}_{0}.$
Let $\mathbf{a}_{i}=(a_{i,1},\ldots,a_{i,n})\in A$ and $\mathbf{a}^{\prime}_{i}=(a^{\prime}_{i,1},\ldots,a^{\prime}_{i,n})\in A$ for all $i\in[1,h]$. We have
$\sum_{i=1}^{h}\mathbf{a}_{i}=\sum_{i=1}^{h}(a_{i,1},\ldots,a_{i,n})=\left(\sum_{i=1}^{h}a_{i,1},\ldots,\sum_{i=1}^{h}a_{i,n}\right)$
and
$\sum_{i=1}^{h}\mathbf{a}^{\prime}_{i}=\sum_{i=1}^{h}(a^{\prime}_{i,1},\ldots,a^{\prime}_{i,n})=\left(\sum_{i=1}^{h}a^{\prime}_{i,1},\ldots,\sum_{i=1}^{h}a^{\prime}_{i,n}\right).$
Moreover,
$f(\mathbf{a}_{i})=\sum_{j=1}^{n}a_{i,j}g^{j-1}\operatorname{\qquad\text{and}\qquad}f(\mathbf{a}^{\prime}_{i})=\sum_{j=1}^{n}a^{\prime}_{i,j}g^{j-1}.$
We have
$\sum_{i=1}^{h}f(\mathbf{a}_{i})=\sum_{i=1}^{h}\sum_{j=1}^{n}a_{i,j}g^{j-1}=\sum_{j=1}^{n}\sum_{i=1}^{h}a_{i,j}g^{j-1}$
and, for all $j\in[1,n]$,
$0\leq\sum_{i=1}^{h}a_{i,j}\leq\sum_{i=1}^{h}\|\mathbf{a}_{i}\|_{\infty}\leq h\max\{\|\mathbf{a}\|_{\infty}:\mathbf{a}\in A\}<g.$
Similarly,
$\sum_{i=1}^{h}f(\mathbf{a}^{\prime}_{i})=\sum_{i=1}^{h}\sum_{j=1}^{n}a^{\prime}_{i,j}g^{j-1}=\sum_{j=1}^{n}\sum_{i=1}^{h}a^{\prime}_{i,j}g^{j-1}$
and, for all $j\in[1,n]$,
$0\leq\sum_{i=1}^{h}a^{\prime}_{i,j}\leq\sum_{i=1}^{h}\|\mathbf{a}^{\prime}_{i}\|_{\infty}\leq h\max\{\|\mathbf{a}^{\prime}\|_{\infty}:\mathbf{a}^{\prime}\in A\}<g.$
It follows from the uniqueness of the $g$-adic representation that
$\sum_{i=1}^{h}f(\mathbf{a}_{i})=\sum_{i=1}^{h}f(\mathbf{a}^{\prime}_{i})$
if and only if, for all $j\in[1,n]$,
$\sum_{i=1}^{h}a_{i,j}=\sum_{i=1}^{h}a^{\prime}_{i,j}$
if and only if
$\sum_{i=1}^{h}\mathbf{a}_{i}=\sum_{i=1}^{h}\mathbf{a}^{\prime}_{i}$
and so the sets $A\subseteq\mathbf{Z}^{n}$ and $B\subseteq\mathbf{Z}$ are Freiman isomorphic of order $h$. This completes the proof. ∎
4. Compression theorem for sums of sets of lattice points
Let $(\mathbf{x},\mathbf{y})$ be an inner product on $\mathbf{R}^{n}$ and let $\mathbf{u}$ be a unit vector in $\mathbf{R}^{n}$. For all $r\in\mathbf{R}$, we have the hyperplane
$H_{\mathbf{u}}(r)=\{\mathbf{x}\in\mathbf{R}^{n}:(\mathbf{u},\mathbf{x})=r\}.$
The $(s,t)$-slab of $\mathbf{u}$ in $\mathbf{R}^{n}$ is the convex set
(14) $H_{\mathbf{u}}(s,t)=\{\mathbf{x}\in\mathbf{R}^{n}:s\leq(\mathbf{u},\mathbf{x})\leq t\}.$
We have $H_{\mathbf{u}}(s,t)=\emptyset$ if $s>t$. If $r<s$, then
(15) $H_{\mathbf{u}}(q,r)\cap H_{\mathbf{u}}(s,t)=\emptyset.$
For $q\leq r$ and $s\leq t$, there is the additive relation
(16) $H_{\mathbf{u}}(q,r)+H_{\mathbf{u}}(s,t)\subseteq H_{\mathbf{u}}(q+s,r+t).$
The diameter with respect to the unit vector $\mathbf{u}$ of a nonempty subset $A$ of $\mathbf{R}^{n}$ is
$\operatorname{\text{diam}}_{\mathbf{u}}(A)=\sup\{(\mathbf{a},\mathbf{u})-(\mathbf{a}^{\prime},\mathbf{u}):\mathbf{a},\mathbf{a}^{\prime}\in A\}.$
Theorem 6.
Let $A$ be a finite subset of $\mathbf{R}^{n}$ with $|A|=k$. Let $\mathbf{u}$ be a unit vector in $\mathbf{R}^{n}$. Enumerate the vectors in $A=\{\mathbf{a}_{1},\mathbf{a}_{2},\ldots,\mathbf{a}_{k}\}$ so that
(17) $(\mathbf{u},\mathbf{a}_{1})\leq(\mathbf{u},\mathbf{a}_{2})\leq\cdots\leq(\mathbf{u},\mathbf{a}_{k}).$
Let $h\geq 2$ and suppose that, for some $j\in[1,k-1]$, there exist $\alpha_{j}>0$ and $\delta_{j}>0$ such that
(18) $(\mathbf{u},\mathbf{a}_{j+1}-\mathbf{a}_{j})=\alpha_{j}+\delta_{j}+(h-1)\max\left((\mathbf{u},\mathbf{a}_{j}-\mathbf{a}_{1}),(\mathbf{u},\mathbf{a}_{k}-\mathbf{a}_{j+1})\right).$
The integer $j$ is unique. There is a subset $A^{\prime}=\{\mathbf{a}^{\prime}_{1},\ldots,\mathbf{a}^{\prime}_{k}\}$ of $\mathbf{R}^{n}$ with $|A^{\prime}|=k$ and
$(\mathbf{u},\mathbf{a}^{\prime}_{1})\leq(\mathbf{u},\mathbf{a}^{\prime}_{2})\leq\cdots\leq(\mathbf{u},\mathbf{a}^{\prime}_{k})$
such that $|hA|=|hA^{\prime}|$ and
(19) $(\mathbf{u},\mathbf{a}^{\prime}_{j+1}-\mathbf{a}^{\prime}_{j})=\alpha_{j}+(h-1)\max((\mathbf{u},\mathbf{a}^{\prime}_{j}-\mathbf{a}^{\prime}_{1}),(\mathbf{u},\mathbf{a}^{\prime}_{k}-\mathbf{a}^{\prime}_{j+1}))$
Moreover,
$\operatorname{\text{diam}}_{\mathbf{u}}(A^{\prime})=\operatorname{\text{diam}}_{\mathbf{u}}(A)-\delta_{j}.$
Proof.
Suppose there exist distinct integers $j_{1}\in[1,k-1]$ and $j_{2}\in[1,k-1]$ that satisfy (18). If $j_{1}<j_{2}$, then
$(\mathbf{u},\mathbf{a}_{j_{1}+1}-\mathbf{a}_{j_{1}})$ $>(h-1)\max((\mathbf{u},\mathbf{a}_{j_{1}}-\mathbf{a}_{1}),(\mathbf{u},\mathbf{a}_{k}-\mathbf{a}_{j_{1}+1}))$
$\geq(h-1)(\mathbf{u},\mathbf{a}_{k}-\mathbf{a}_{j_{1}+1})$
$\geq(\mathbf{u},\mathbf{a}_{j_{2}+1}-\mathbf{a}_{j_{2}})$
$>(h-1)\max((\mathbf{u},\mathbf{a}_{j_{2}}-\mathbf{a}_{1}),(\mathbf{u},\mathbf{a}_{k}-\mathbf{a}_{j_{2}+1}))$
$\geq(h-1)(\mathbf{u},\mathbf{a}_{j_{2}}-\mathbf{a}_{1})$
$\geq(\mathbf{u},\mathbf{a}_{j_{1}+1}-\mathbf{a}_{j_{1}})$
which is absurd. Therefore, at most one integer $j\in[1,k-1]$ satisfies (18).
Suppose that relation (18) holds for some $j\in[1,h-1]$. Let
$q=(\mathbf{u},\mathbf{a}_{1})\qquad$ $\qquad r=(\mathbf{u},\mathbf{a}_{j})$
$s=(\mathbf{u},\mathbf{a}_{j+1})\qquad$ $\qquad t=(\mathbf{u},\mathbf{a}_{k})$
Then
$q\leq r<s\leq t$
and
(20) $s-r=\alpha_{j}+\delta_{j}+(h-1)\max(r-q,t-s).$
Let
$A=B\cup C$
where
$B=\{\mathbf{a}_{1},\ldots,\mathbf{a}_{j}\}\subseteq H_{\mathbf{u}}(q,r)$
and
$C=\{\mathbf{a}_{j+1},\ldots,\mathbf{a}_{k}\}\subseteq H_{\mathbf{u}}(s,t)$
and $H_{\mathbf{u}}(q,r)$ and $H_{\mathbf{u}}(s,t)$ are the slabs defined by (14). We have
$hA=\bigcup_{i=0}^{h}\left((h-i)B+iC\right).$
The additive relation (16) implies that
$(h-i)B\subseteq H_{\mathbf{u}}((h-i)q,(h-i)r)$
and
$iC\subseteq H_{\mathbf{u}}(is,it)$
and so
$(h-i)B+iC\subseteq H_{\mathbf{u}}((h-i)q+is,(h-i)r+it).$
Let $i\in[0,h-1]$. If
(hi)r+it<(hi1)q+(i+1)s(h-i)r+it<(h-i-1)q+(i+1)s
or, equivalently, if
(h1i)(rq)+i(ts)<sr(h-1-i)(r-q)+i(t-s)<s-r
then, by (15), the sets (hi)B+iC(h-i)B+iC and (h1i)B+(i+1)C(h-1-i)B+(i+1)C are disjoint.
From (20) we have
(21) $(h-1-i)(r-q)+i(t-s)\leq(h-1)\max(r-q,t-s)<s-r$
and so the sets (hi)B+iC(h-i)B+iC are pairwise disjoint for all $i\in[0,h]$. It follows that
$|hA|=\sum_{i=0}^{h}\left|(h-i)B+iC\right|.$
Let $\mathbf{a}^{\prime}_{\ell}=\mathbf{a}_{\ell}$ for all $\ell\in[1,j]$ and let
$B^{\prime}=\{\mathbf{a}^{\prime}_{1},\ldots,\mathbf{a}^{\prime}_{j}\}=B.$
Then
$B^{\prime}\subseteq H_{\mathbf{u}}(q,r).$
Let $\mathbf{a}^{\prime}_{\ell}=\mathbf{a}_{\ell}-\delta_{j}\mathbf{u}$ for all $\ell\in[j+1,k]$ and let
$C^{\prime}=\{\mathbf{a}^{\prime}_{j+1},\ldots,\mathbf{a}^{\prime}_{k}\}=C-\delta_{j}\mathbf{u}.$
Because $\mathbf{u}$ is a unit vector, we have $\delta_{j}=(\mathbf{u},\delta_{j}\mathbf{u})$. For all $\ell\in[j+1,k]$, we have
$s-\delta_{j}$ $=(\mathbf{u},\mathbf{a}_{j+1})-\delta_{j}=(\mathbf{u},\mathbf{a}_{j+1})-(\mathbf{u},\ \delta_{j}\mathbf{u})$
$=(\mathbf{u},\mathbf{a}_{j+1}-\delta_{j}\mathbf{u})=(\mathbf{u},\mathbf{a}^{\prime}_{j+1})$
$\leq(\mathbf{u},\mathbf{a}_{\ell})-\delta_{j}=(\mathbf{u},\mathbf{a}^{\prime}_{\ell})$
$\leq(\mathbf{u},\mathbf{a}_{k})-\delta_{j}=(\mathbf{u},\mathbf{a}^{\prime}_{k})$
$=t-\delta_{j}$
and so
$C^{\prime}\subseteq H_{\mathbf{u}}(s-\delta_{j},t-\delta_{j}).$
Let $A^{\prime}=B^{\prime}\cup C^{\prime}$. We have
$hA^{\prime}=\bigcup_{i=0}^{h}\left((h-i)B^{\prime}+iC^{\prime}\right)$
and
$(h-i)B^{\prime}+iC^{\prime}\subseteq H_{\mathbf{u}}((h-i)q+i(s-\delta_{j}),(h-i)r+i(t-\delta_{j}))$
for all $i\in[0,h]$. The $h+1$ sumsets $(h-i)B^{\prime}+iC^{\prime}$ are pairwise disjoint if
$(h-i)r+i(t-\delta_{j})<(h-1-i)q+(i+1)(s-\delta_{j})$
or, equivalently, if
$(h-1-i)(r-q)+i(t-s)<s-r-\delta_{j}$
for all $i\in[0,h-1]$. From relation (18), we have
$(h-1-i)$ $(r-q)+i(t-s)\leq(h-1)\max(r-q,t-s)$
$=(h-1)\max\left((\mathbf{u},\mathbf{a}_{j}-\mathbf{a}_{1}),(\mathbf{u},\mathbf{a}_{k}-\mathbf{a}_{j+1})\right)$
$=(\mathbf{u},\mathbf{a}_{j+1}-\mathbf{a}_{j})-\alpha_{j}-\delta_{j}$
$<s-r-\delta_{j}.$
Thus, the sets $(h-i)B^{\prime}+iC^{\prime}$ pairwise disjoint for $i\in[0,h]$ and
$|hA^{\prime}|=\sum_{i=1}^{h}|(h-i)B^{\prime}+iC^{\prime}|.$
Because $B^{\prime}=B$ and $C^{\prime}=C-\delta_{j}\mathbf{u}$, we have
$(h-i)B^{\prime}+iC^{\prime}=(h-i)B+iC-i\delta_{j}\mathbf{u}$
and so
$|(h-i)B^{\prime}+iC^{\prime}|=|(h-i)B+iC-i\delta_{j}\mathbf{u}|=|(h-i)B+iC|$
and
$|hA|=\sum_{i=0}^{h}|(h-i)B+iC|=\sum_{i=0}^{h}|(h-i)B^{\prime}+iC^{\prime}|=|hA^{\prime}|.$
Moreover,
$\operatorname{\text{diam}}_{\mathbf{u}}(A^{\prime})=(\mathbf{u},\mathbf{a}^{\prime}_{k})-(\mathbf{u},\mathbf{a}^{\prime}_{1})=(\mathbf{u},\mathbf{a}_{k})-(\mathbf{u},\mathbf{a}_{1})-\delta_{j}=\operatorname{\text{diam}}_{\mathbf{u}}(A)-\delta_{j}.$
This completes the proof. ∎
Theorem 7.
Let $A=\{\mathbf{a}_{1},\ldots,\mathbf{a}_{k}\}$ be a finite subset of $\mathbf{Z}^{n}$ with $|A|=k$. Let $\mathbf{e}_{1},\ldots,\mathbf{e}_{n}$ be the standard basis for $\mathbf{R}^{n}$. For all $p\in[1,n]$, there is a permutation $\sigma_{p}$ of the integer interval $[1,k]$ such that
(22) $(\mathbf{e}_{p},\mathbf{a}_{\sigma_{p}(1)})\leq(\mathbf{e}_{p},\mathbf{a}_{\sigma_{p}(2)})\leq\cdots\leq(\mathbf{e}_{p},\mathbf{a}_{\sigma_{p}(k)}).$
There exists a set $A^{*}=\{\mathbf{a}^{*}_{1},\ldots,\mathbf{a}^{*}_{k}\}$ in $\mathbf{Z}^{n}$ with $\left|A^{*}\right|=k$ such that $|hA|=|hA^{*}|$ and, for all $p\in[1,n]$ and $j\in[1,k-1]$,
$\left(\mathbf{e}_{p},\mathbf{a}^{*}_{\sigma_{p}(j+1)}-\mathbf{a}^{*}_{\sigma_{p}(j)}\right)$
$\qquad\leq 1+(h-1)\max\left(\left(\mathbf{e}_{p},\mathbf{a}^{*}_{\sigma_{p}(j)}-\mathbf{a}^{*}_{\sigma_{p}(1)}\right),\left(\mathbf{e}_{p},\mathbf{a}^{*}_{\sigma_{p}(k)}-\mathbf{a}^{*}_{\sigma_{p}(j+1)}\right)\right).$
Proof.
Because the standard basis vectors $\mathbf{e}_{1},\ldots,\mathbf{e}_{n}$ are pairwise orthogonal, the monotonicity inequalities (22) persist if each vector $\mathbf{a}_{j}$ is replaced by a vector of the form $\mathbf{a}_{j}-\mathbf{v}_{j}$ for any vector $\mathbf{v}_{j}$ in the hyperplane $H_{\mathbf{e}_{p}}(0)$, that is, by any vector that is a linear combination of the n1n-1 vectors in the set $\{\mathbf{e}_{1},\ldots,\mathbf{e}_{n}\}\setminus\{\mathbf{e}_{p}\}$.
If
$\left(\mathbf{e}_{1},\mathbf{a}_{\sigma_{1}(j+1)}-\mathbf{a}_{\sigma_{1}(j)}\right)$
$\qquad>1+(h-1)\max\left(\left(\mathbf{e}_{1},\mathbf{a}_{\sigma_{1}(j)}-\mathbf{a}_{\sigma_{1}(1)}\right),\left(\mathbf{e}_{1},\mathbf{a}_{\sigma_{1}(k)}-\mathbf{a}_{\sigma_{1}(j+1)}\right)\right)$
for some $j\in[1,k-1]$, then there is a positive integer $\delta_{j}$ such that
$\left(\mathbf{e}_{1},\mathbf{a}_{\sigma_{1}(j+1)}-\mathbf{a}_{\sigma_{1}(j)}\right)$
$\qquad=1+\delta_{j}+(h-1)\max\left(\left(\mathbf{e}_{1},\mathbf{a}_{\sigma_{1}(j)}-\mathbf{a}_{\sigma_{1}(1)}\right),\left(\mathbf{e}_{1},\mathbf{a}_{\sigma_{1}(k)}-\mathbf{a}_{\sigma_{1}(j+1)}\right)\right).$
Applying Theorem 6 with the unit vector $\mathbf{e}_{1}$ and $\alpha_{j}=1$, we obtain a set $A^{\prime}=\{\mathbf{a}^{\prime}_{1},\ldots,\mathbf{a}^{\prime}_{k}\}$ in $\mathbf{Z}^{n}$ with $|A^{\prime}|=k$ such that $|hA|=|hA^{\prime}|$ and
$\operatorname{\text{diam}}_{\mathbf{e}_{1}}(A^{\prime})=\operatorname{\text{diam}}_{\mathbf{e}_{1}}(A)-\delta_{j}\leq\operatorname{\text{diam}}_{\mathbf{e}_{1}}(A)-1.$
Moreover, for all $j\in[1,k]$, there is a scalar $c_{1,j}$ such that $\mathbf{a}^{\prime}_{j}=\mathbf{a}_{j}-c_{1,j}\mathbf{e}_{1}$. If
$\left(\mathbf{e}_{1},\mathbf{a}^{\prime}_{\sigma_{1}(j+1)}-\mathbf{a}^{\prime}_{\sigma_{1}(j)}\right)$
$\qquad>1+(h-1)\max\left(\left(\mathbf{e}_{1},\mathbf{a}^{\prime}_{\sigma_{1}(j)}-\mathbf{a}^{\prime}_{\sigma_{1}(1)}\right),\left(\mathbf{e}_{1},\mathbf{a}^{\prime}_{\sigma_{1}(k)}-\mathbf{a}^{\prime}_{\sigma_{1}(j+1)}\right)\right)$
for some $j^{\prime}\in[1,k-1]$, then we again apply Theorem 6 obtain a set $A^{\prime\prime}$ in $\mathbf{Z}^{n}$ with $|A^{\prime\prime}|=k$ such that $|hA|=|hA^{\prime\prime}|$ and
$\operatorname{\text{diam}}_{\mathbf{e}_{1}}(A^{\prime\prime})\leq\operatorname{\text{diam}}_{\mathbf{e}_{1}}(A^{\prime})-1\leq\operatorname{\text{diam}}_{\mathbf{e}_{1}}(A)-2.$
Iterating this process, we obtain a sequence of sets whose diameters form a strictly decreasing sequence of positive integers. This process must terminate after finitely many iterations and we obtain a set $A^{(1)}=\left\{\mathbf{a}_{1}^{(1)},\ldots,\mathbf{a}^{(1)}_{k}\right\}$ with $\left|A^{(1)}\right|=k$ such that $|hA^{(1)}|=|hA|$ and
$\left(\mathbf{e}_{1},\mathbf{a}^{\prime}_{\sigma_{1}(j+1)}-\mathbf{a}^{\prime}_{\sigma_{1}(j)}\right)$
$\qquad\leq 1+(h-1)\max\left(\left(\mathbf{e}_{1},\mathbf{a}^{\prime}_{\sigma_{1}(j)}-\mathbf{a}^{\prime}_{\sigma_{1}(1)}\right),\left(\mathbf{e}_{1},\mathbf{a}^{\prime}_{\sigma_{1}(k)}-\mathbf{a}^{\prime}_{\sigma_{1}(j+1)}\right)\right)$
for all $j\in[1,k-1]$. Moreover, each vector $\mathbf{a}_{j}^{(1)}$ in $A^{(1)}$ differs from the vector $\mathbf{a}_{j}$ in $A$ by a scalar multiple of $\mathbf{e}_{1}$.
Suppose for some $p\in[1,n-1]$ we have constructed a set $A^{(p)}$ in $\mathbf{Z}^{n}$ with $\left|A^{(p)}\right|=k$ such that $|hA^{(p)}|=|hA|$ and, for all $p^{\prime}\in[1,p]$ and for all $j\in[1,k-1]$,
$\left(\mathbf{e}_{p^{\prime}},\mathbf{a}^{\left(p\right)}_{\sigma_{p^{\prime}}(j+1)}-\mathbf{a}^{\left(p\right)}_{\sigma_{p^{\prime}}(j)}\right)$
$\qquad\leq 1+\left(h-1\right)\max\left(\left(\mathbf{e}_{p^{\prime}},\mathbf{a}^{\left(p\right)}_{\sigma_{p^{\prime}}(j)}-\mathbf{a}^{\left(p\right)}_{\sigma_{p^{\prime}}(1)}\right),\left(\mathbf{e}_{p^{\prime}},\mathbf{a}^{\left(p\right)}_{\sigma_{p^{\prime}}(k)}-\mathbf{a}^{\left(p\right)}_{\sigma_{p^{\prime}}(j+1)}\right)\right).$
Moreover, each vector in $A^{(p)}$ differs from a vector in $A$ by a linear combination of $\mathbf{e}_{1},\ldots,\mathbf{e}_{p}$.
We apply Theorem 6 to the set $A^{(p)}$ with the unit vector $\mathbf{e}_{p+1}$ and obtain a set $A^{(p+1)}$ with $\left|A^{(p+1)}\right|=k$ such that $\left|hA^{(p+1)}\right|=\left|hA\right|$ and
$\left(\mathbf{e}_{p+1},\mathbf{a}^{\left(p+1\right)}_{j+1}-\mathbf{a}^{\left(p+1\right)}_{j}\right)$
$\quad\leq 1+\left(h-1\right)\max\left(\left(\mathbf{e}_{p+1},\mathbf{a}^{\left(p+1\right)}_{j}-\mathbf{a}^{\left(p+1\right)}_{1}\right),\left(\mathbf{e}_{p+1},\mathbf{a}^{\left(p+1\right)}_{k}-\mathbf{a}^{\left(1\right)}_{j+1}\right)\right)$
for all $j\in[1,k-1]$. Because each vector in $A^{(p+1)}$ differs from a vector in $A$ by a linear combination of $\mathbf{e}_{1},\ldots,\mathbf{e}_{p+1}$, we also have, for all $p^{\prime}\in[1,p]$ and $j\in[1,k-1]$,
$\left(\mathbf{e}_{p^{\prime}},\mathbf{a}^{(p+1)}_{\sigma_{p^{\prime}}(j+1)}-\mathbf{a}^{(p+1)}_{\sigma_{p^{\prime}}(j)}\right)$
$\qquad\leq 1+\left(h-1\right)\max\left(\left(\mathbf{e}_{p^{\prime}},\mathbf{a}^{(p+1)}_{\sigma_{p^{\prime}}(j)}-\mathbf{a}^{(p+1)}_{\sigma_{p^{\prime}}(1)}\right),\left(\mathbf{e}_{p^{\prime}},\mathbf{a}^{(p+1)}_{\sigma_{p^{\prime}}(k)}-\mathbf{a}^{(p+1)}_{\sigma_{p^{\prime}}(j+1)}\right)\right).$
Continuing inductively completes the proof. ∎
5. An upper bound for $N(h,k)$
The computational complexity of the sumset size problem is determined by the integer $N(h,k)$, which is the smallest integer $N$ such that
$\mathcal{R}_{\mathbf{Z}}(h,k)=\{|hA|:A\subseteq[0,N-1]\text{ and }|A|=k\}.$
In this section we compute an upper bound for $N(h,k)$. This result is based on work with Noah Kravitz.
Let $p$ be a prime number and let $(\mathbf{Z}/p\mathbf{Z})^{*}$ be the multiplicative group of nonzero congruence classes mod $p$. For all $a\in\mathbf{Z}$, let $\overline{a}$ denote the congruence class $a+p\mathbf{Z}$. For $\overline{A}\subseteq\mathbf{Z}/p\mathbf{Z}$ and $\lambda\in(\mathbf{Z}/p\mathbf{Z})^{*}$, let
$\lambda\ast\overline{A}=\{\lambda\overline{a}:\overline{a}\in\overline{A}\}.$
The following result is known. A proof is included for completeness.
Lemma 1.
(a) For $\overline{A}\subseteq\mathbf{Z}/p\mathbf{Z}$ and $\lambda\in(\mathbf{Z}/p\mathbf{Z})^{*}$, the sets $\overline{A}$ and $\lambda\ast\overline{A}$ are Freiman isomorphic of order $h$ for all $h\geq 2$.
(b) Let $A=\{a_{1},\ldots,a_{k}\}$ be a set of $k$ integers and let $\overline{A}=\{\overline{a_{1}},\ldots,\overline{a_{k}}\}$ be the corresponding set of congruence classes in the field $\mathbf{Z}/p\mathbf{Z}$. Let $h\geq 2$. If $|a_{i}|<p/2h$ for all $i\in[1,k]$, then the sets $A$ and $\overline{A}$ are Freiman isomorphic of order $h$.
Proof.
To prove (a), it suffices to observe that, for all $\lambda\in(\mathbf{Z}/p\mathbf{Z})^{*}$, the function $f:\overline{A}\rightarrow\lambda\ast\overline{A}$ defined by $f(\overline{a})=\lambda\overline{a}$ is a Freiman isomorphism of order $h$ for all $h\geq 2$.
To prove (b), consider the function $g:A\rightarrow\overline{A}$ defined by $g(a)=\overline{a}$. Because $|a_{i}|<p/2h<p/2$ for all $i\in[1,k]$, the function $g$ is a bijection. Let $a_{i},b_{i}\in A$ for all $i\in[1,h]$. If
$\sum_{i=1}^{h}a_{i}=\sum_{i=1}^{h}b_{i}$
then
$\sum_{i=1}^{h}g(a_{i})=\sum_{i=1}^{h}\overline{a_{i}}=\sum_{i=1}^{h}a_{i}+p\mathbf{Z}=\sum_{i=1}^{h}b_{i}+p\mathbf{Z}=\sum_{i=1}^{h}\overline{b_{i}}=\sum_{i=1}^{h}g(b_{i}).$
Conversely, if
$\sum_{i=1}^{h}g(a_{i})=\sum_{i=1}^{h}g(b_{i})$
then
$\sum_{i=1}^{h}a_{i}+p\mathbf{Z}=\sum_{i=1}^{h}\overline{a_{i}}=\sum_{i=1}^{h}\overline{b_{i}}=\sum_{i=1}^{h}b_{i}+p\mathbf{Z}$
and so $p$ divides $\sum_{i=1}^{h}b_{i}-\sum_{i=1}^{h}a_{i}$. Because $|a_{i}|<p/2h$ for all $i\in[1,k]$, we have
$\left|\sum_{i=1}^{h}a_{i}\right|\leq\sum_{i=1}^{h}|a_{i}|<\frac{p}{2}\operatorname{\qquad\text{and}\qquad}\left|\sum_{i=1}^{h}b_{i}\right|\leq\sum_{i=1}^{h}|b_{i}|<\frac{p}{2}$
and so
$p<\sum_{i=1}^{h}b_{i}-\sum_{i=1}^{h}a_{i}<p.$
It follows that $\sum_{i=1}^{h}a_{i}=\sum_{i=1}^{h}b_{i}$, and so $g$ is an Freiman isomorphism of order $h$. This completes the proof. ∎
Let $p$ be an odd prime. For $\overline{x}=x+p\mathbf{Z}\in\mathbf{Z}/p\mathbf{Z}$, let $|\overline{x}|$ be the smallest absolute value of an integer in the congruence class $\overline{x}$. Thus, $|\overline{x}|\in[0,(p-1)/2]$. For $\mathbf{x}=(\overline{x_{1}},\ldots,\overline{x_{k}})\in(\mathbf{Z}/p\mathbf{Z})^{k}$, let
$\|\mathbf{x}\|_{\infty}=\max\{|\overline{x_{j}}|:j\in[1,k]\}\in\left[0,\frac{p-1}{2}\right].$
For all $\mathbf{v}_{1},\mathbf{v}_{2}\in(\mathbf{Z}/p\mathbf{Z})^{k}$, we have
(23) $\|\mathbf{v}_{1}\pm\mathbf{v}_{2}\|_{\infty}\leq\|\mathbf{v}_{1}\|_{\infty}+\|\mathbf{v}_{2}\|_{\infty}.$
Let $\mathbf{0}=(\overline{0},\ldots,\overline{0})\in(\mathbf{Z}/p\mathbf{Z})^{k}$. For $\mathbf{x}\in(\mathbf{Z}/p\mathbf{Z})^{k}$ and $r\in[0,(p-1)/2]$, define the boxes
$B_{r}(\mathbf{0})=\left\{\mathbf{v}\in(\mathbf{Z}/p\mathbf{Z})^{k}:\|\mathbf{v}\|_{\infty}\leq r\right\}$
and
$B_{r}(\mathbf{x})=\left\{\mathbf{x}+\mathbf{v}:\mathbf{v}\in B_{r}(\mathbf{0})\right\}=\mathbf{x}+B_{r}(\mathbf{0}).$
We have
(24) $\left|B_{r}(\mathbf{0})\right|=\left|B_{r}(\mathbf{x})\right|=(2r+1)^{k}.$
Theorem 8.
For all $h\geq 3$ and $k\geq 3$,
$N(h,k)<4(8h)^{k-1}.$
Proof.
Let $t\in\mathcal{R}_{\mathbf{Z}}(h,k)$ and let $M=M(t)$ be the smallest positive integer such that there exists a set
$A=\{a_{1},\ldots,a_{k}\}\subseteq[-M,M]$
with $|A|=k$ and $|hA|=t$. We shall prove that $M(t)<2(8h)^{k-1}$ for all $t\in\mathcal{R}_{\mathbf{Z}}(h,k)$. It follows that $N(h,k)<4(8h)^{k-1}$.
By Bertrand’s postulate, there is a prime $p$ such that
(25) $2hM<p<4hM$
and so
$|a_{j}|\leq M<\frac{p}{2h}$
for all $j\in[1,k]$. Let
$\overline{A}=\{\overline{a_{1}},\ldots,\overline{a_{k}}\}\subseteq\mathbf{Z}/p\mathbf{Z}.$
By Lemma 1(b), the sets $A$ and $\overline{A}$ are Freiman isomorphic of order $h$. By Lemma 1(a), the sets $\overline{A}$ and $\lambda\ast\overline{A}$ are Freiman isomorphic of order $h$ for all $\lambda\in(\mathbf{Z}/p\mathbf{Z})^{*}$.
Choose the integer $r$ such that
$\frac{1}{2}\left(p^{1-(1/k)}-1\right)<r\leq\frac{1}{2}\left(p^{1-(1/k)}+1\right).$
Then
(26) $p(2r+1)^{k}>p^{k}.$
We may assume that $M\geq 2(4h)^{k-1}$ and so
$(4h)^{k}\leq 2hM<p.$
It follows that
(27) $2r\leq p^{1-(1/k)}+1<\frac{2p}{p^{1/k}}<\frac{p}{2h}.$
Fix an ordering of the $k$ congruence classes in the set $\overline{A}$ and consider the vector
$\mathbf{a}=(\overline{a_{1}},\ldots,\overline{a_{k}})\in(\mathbf{Z}/p\mathbf{Z})^{k}.$
For all $\lambda\in\mathbf{Z}/p\mathbf{Z}$, we have the vector
$\lambda\mathbf{a}=(\lambda\overline{a_{1}},\ldots,\lambda\overline{a_{k}})$
and the box $B_{r}(\lambda\mathbf{a})$ in $(\mathbf{Z}/p\mathbf{Z})^{k}$. From relations (24) and (26), we obtain
$\sum_{\lambda\in\mathbf{Z}/p\mathbf{Z}}\left|B_{r}(\lambda\mathbf{a})\right|=p(2r+1)^{k}>p^{k}=\left|(\mathbf{Z}/p\mathbf{Z})^{k}\right|$
and so the $p$ boxes $B_{r}(\lambda\mathbf{a})$ are not pairwise disjoint. It follows that there exist $\lambda_{1},\lambda_{2}\in\mathbf{Z}/p\mathbf{Z}$ such that $\lambda_{1}\neq\lambda_{2}$ and
$B_{r}(\lambda_{1}\mathbf{a})\cap B_{r}(\lambda_{2}\mathbf{a})\neq\emptyset.$
We obtain $\mathbf{v}_{1},\mathbf{v}_{2}\in B_{r}(\mathbf{0})$ such that
$\lambda_{1}\mathbf{a}+\mathbf{v}_{1}=\lambda_{2}\mathbf{a}+\mathbf{v}_{2}.$
Let
$\lambda=\lambda_{1}-\lambda_{2}\in(\mathbf{Z}/p\mathbf{Z})^{*}$
and
$\mathbf{b}=\lambda\mathbf{a}=\lambda_{1}\mathbf{a}-\lambda_{2}\mathbf{a}=\mathbf{v}_{2}-\mathbf{v}_{1}\in(\mathbf{Z}/p\mathbf{Z})^{k}.$
From inequalities (23) and  (27), we have
$\|\mathbf{b}\|_{\infty}=\|\mathbf{v}_{2}-\mathbf{v}_{1}\|_{\infty}\leq\|\mathbf{v}_{1}\|_{\infty}+\|\mathbf{v}_{2}\|_{\infty}\leq 2r<\frac{p}{2h}$
and so there are integers $b_{1},\ldots,b_{k}$ such that
$\mathbf{b}=(\overline{b_{1}},\ldots,\overline{b_{k}})=\lambda\mathbf{a}$
and
$|b_{j}|\leq 2r<\frac{p}{2h}$
for all $j\in[1,h]$. Let
$\overline{B}=\{\overline{b_{1}},\ldots,\overline{b_{k}}\}=\lambda\ast\overline{A}.$
By Lemma 1, the sets of congruence classes $\overline{A}$ and $\overline{B}$ are Freiman isomorphic of order $h$, and the set of congruence classes $\overline{B}$ and the set of integers $B=\{b_{1},\ldots,b_{k}\}$ are Freiman isomorphic of order $h$. Therefore, $|hB|=|h\overline{B}|=|h\overline{A}|=|hA|=t$. Moreover,
$B\subseteq[-2r,2r].$
The minimality of $M$ implies
$M$ $\leq 2r\leq 2p^{1-(1/k)}<2(4hM)^{1-(1/k)}.$
Solving for $M$, we obtain
$M<2(8h)^{k-1}.$
This completes the proof. ∎